If Jim's boat goes 30 mph upstream and 50 mph downsteam. What is the speed of the current and the boat in still water? I HAVE NO IDEA HOW TO SET THIS UP!! thanks!
Answer by ankor@dixie-net.com(22740)

If Jim's boat goes 30 mph upstream and 50 mph downsteam.
What is the speed of the current and the boat in still water?
:
Let s = speed of the boat in still water
Let c = speed of the current
:
s + c = 50; downstream
s - c = 30; upstream
-------------------------adding eliminates c find s
2s = 80
s = 80/2
s = 40 mph in still water
:
Find the current using the 1st equation
40 + c = 50
c = 50 - 40
c = 10 mph is current